### v3

# and application to tori over local fields that split over cyclic extensions

Abstract.
In this note we give formulas for cup product in Tate cohomology in terms of inhomogeneous cochains. Using one of these formulas, for a torus $T$ defined over a non-archimedean local field $K$ and splitting over a cyclic extension of $K$, we compute explicit cocycles representing all cohomology classes in $H^{1}(K,T)$.
Key words and phrases:
Galois cohomology, local fields, local class field theory, algebraic tori, Tate cohomology, cup product
1991 Mathematics Subject Classification:
11E72, 11R37 , 11S , 20G10, 20G25, 20J06
This research was partially supported by the Israel Science Foundation (grant 1030/22).
1. Introduction
$1.$ The homogeneous and inhomogeneous complexes
Let $G$ be a finite group and AA be a $G$-module. We consider the homogeneous cochain complex $F^{\scriptscriptstyle\bullet}$ for AA (see [2, Appendix E]):
(1) ${\dots F^{-2}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{-1}}$${F^{-1}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{0}}$${F^{0}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{1}}$${F^{1}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{2}}$${F^{2}(G,A)\dots}$${{\rm Maps}_{G}(G^{2},A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm Maps}_{G}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm Maps}_{G}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm Maps}_{G}(G^{2},A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm Maps}_{G}(G^{3},A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
We also consider the inhomogeneous cochain complex $C^{\scriptscriptstyle\bullet}$ for AA (see [2, Appendix E]):
(2) ${\dots C^{-2}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{-1}}$${C^{-1}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{0}}$${C^{0}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{1}}$${C^{1}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d^{2}}$${C^{2}(G,A)\dots}$${{\rm Maps}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm Maps}(G,A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm Maps}(G^{2},A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
These two complexes $F^{\scriptscriptstyle\bullet}$ and $C^{\scriptscriptstyle\bullet}$ are canonically isomorphic; see Section 2 below. For $p\in{\mathds{Z}}$, the elements of $F^{p}(G,A)$ are called homogeneous $p$-cochains, and the elements of $C^{p}(G,A)$ are called inhomogeneous $p$-cochains.
For $p,q\in{\mathds{Z}}$, there is a canonical cup product pairing
$\cup\colon F^{p}(G,A)\times F^{q}(G,A)\to F^{p+q}(G,A);$
see [3, Section IV.7] or Section 2 below. In Section 2, using the canonical isomorphism between the homogeneous and the inhomogeneous complexes, we give formulas for the cup product pairing in terms of inhomogeneous cochains
$\cup\colon C^{p}(G,A)\times C^{q}(G,A)\to C^{p+q}(G,A).$
1.2. Tori over local fields
Let $K$ be a non-archimedean local field, and $T$ be a $K$-torus. Let $L/K$ be a finite Galois extension splitting $T$. Consider the Galois group $G={\rm Gal}(L/K)$ and the cocharacter group
$X={\sf X}_{*}(T):={\rm Hom}({\mathds{G}}_{m,L},T_{L}).$
Then $G$ naturally acts on $X$. Consider the first Galois cohomology group
$H^{1}(K,T)=H^{1}(L/K,T)\coloneqq H^{1}(G,T(L))=H^{1}(G,X\otimes L^{\times}).$
According to Tate [7], there is a canonical isomorphism
(3) $H^{-1}(G,X){\ \xrightarrow{\makebox[9.56667pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }H^{1}(G,X\otimes L^{\times}).$
Recall that, denoting by $X_{G}$ the group of coinvariants of $G$ in $X$:
$X_{G}=X/\langle x-g\cdot x\ |\ g\in G,x\in X\rangle,$
the group $H^{-1}(G,X)$ is the subgroup of $X_{G}$ consisting of the classes of elements $x\in X$ with $\sum_{g\in G}\,g\kern-0.8pt\cdot\kern-0.8ptx=0$. Let
$u_{L/K}\in H^{2}(G,L^{\times})$
denote the fundamental class; see, for instance, [4, Definition 9.1]. The isomorphism (3) is given by cup product with $u_{L/K}$.
For applications, in particular, for twisting, it would be helpful to have an explicit formula for the isomorphism (3). In the general case this seems to be out of reach. We give a desired formula in the case of a cyclic extension $L/K$, that is, when $G$ is cyclic.
In the case where the extension $L/K$ is unramified, and hence cyclic (that is, $G$ is cyclic), there is a formula for the fundamental class $u_{L/K}$ in Milne’s notes [5]. In the general case of finite Galois extension $L/K$ of local fields, the fundamental class is mysterious. In the answer [6] to a MathOverflow question of the author, Will Sawin guessed and partly proved a formula for $u_{L/K}$ in the case when $L/K$ is cyclic (possibly ramified). We prove Sawin’s formula in full using Milne’s answer to the author’s question [1]:
Theorem 1.1 (Sawin [6]).
Let $L/K$ be a cyclic extension of degree $n$ of non-archimedean local fields. Choose a generator $\sigma$ of the cyclic group (of order $n$) $G={\rm Gal}(L/K)$. Let $e_{\sigma}\in K^{\times}$ be an element such that
$(e_{\sigma},L/K)=\sigma$
where
$e\mapsto(e,L/K)\colon\ K^{\times}\to K^{\times}/{\rm Nm}(L^{\times}){\ \xrightarrow{\makebox[10.13342pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }G$
is the local reciprocity homomorphism; see [3, Section VI.2.2]. Then the fundamental class $u_{L/K}\in H^{2}(G,L^{\times})$ is represented by the following cocycle $a\colon G\times G\to L^{\times}$:
$\text{for \ $0\leq i,j<n$,}\qquad a(\sigma^{i},\sigma^{j})=\begin{cases}1&\text{if \ \ }i+j<n,\\ e_{\sigma}&\text{if \ \ }i+j\geq n.\end{cases}$
Using Theorem 1.1 and one of our formulas for cup product in terms of inhomogeneous cochains, we explicitly compute the isomorphism (3) in the cyclic case:
Theorem 1.2.
Let $T$ be a torus over a non-archimedean local field $K$, and assume that $T$ splits over a cyclic extension $L$ of $K$ of degree $n$. Let $G={\rm Gal}(L/K)$, and let $\sigma\in G$ and $e_{\sigma}\in K^{\times}$ be as in Theorem $1.$ Let $x\in X={\sf X}_{*}(T)$ be a cocharacter such that $x]\in X_{G}$ is contained in $H^{-1}(G,X)$. Then the image of $x]\in H^{-1}(G,X)$ in $H^{1}(G,X\otimes L^{\times})$ under the isomorphism (3) is the class of the cocycle
$z_{x}\colon\,G\to X\,\otimes\,L^{\times},\quad\ \sigma^{g}\,\longmapsto\ \sum_{t=1}^{g}\sigma^{t}\kern-0.8pt\cdot\kern-0.8ptx\,\otimes\,e_{\sigma}\quad\text{for}\ \,0\leq g<n.$
Note that the condition $0\leq g<n$ in the theorem can be replaced by $g\geq 0$.
The plan of the note is as follows. In Section 2 we write the formulas for cup product in terms of homogeneous and inhomogeneous cocycles. In Section 3 we prove Theorem $1.$ In Section 4 we prove Theorem 1.2. In Appendix A we provide details of calculations giving the “inhomogeneous” formulas of Section 2.
Acknowledgements: The author is grateful to Will Sawin for his answer [6] to the author’s MathOverflow question [1], and to Tasho Kaletha and James S. Milne for helpful email correspondence.
2. Cup product in terms of inhomogeneous cocycles
In this section, $G$ is an arbitrary finite group and AA is a $G$-module.
We consider the homogeneous cochain complex (1) for AA, which we denote by $F^{\scriptscriptstyle\bullet}$. We also consider the inhomogeneous complex (2) for AA, which we denote by $C^{\scriptscriptstyle\bullet}$. For details, see [2, Appendix E].
These two complexes $F^{\scriptscriptstyle\bullet}$ and $C^{\scriptscriptstyle\bullet}$ are canonically isomorphic, with the isomorphism given as follows. For a homogeneous cochain $f\in F^{n}(G,A)$ with corresponding inhomogeneous cochain $c\in C^{n}(G,A)$, we have
(4) $c(h_{1},\dots,h_{n})$ $=f(\kern 0.8pt1,\kern 0.8pth_{1},\kern 0.8pth_{1}h_{2},\kern 0.8pt\dots,\kern 0.8pth_{1}\cdots h_{n}\kern 0.8pt),\quad n\geq 0$
(5) $f(g_{0},\dots,g_{n})$ $=g_{0}c(\kern 0.8ptg_{0}^{-1}g_{1},\kern 0.8ptg_{1}^{-1}g_{2},\kern 0.8ptg_{2}^{-1}g_{3},\kern 0.8pt\dots,\kern 0.8ptg_{n-1}^{-1}g_{n}\kern 0.8pt),\quad\ n\geq 0$
(6) $c(h_{1},\dots,h_{-n-1})$ $=f(\kern 0.8pt1,h_{1},\kern 0.8pth_{1}h_{2},\kern 0.8pt\dots,\kern 0.8pth_{1}\cdots h_{-n-1}\kern 0.8pt),\quad n<0$
(7) $f(g_{1},\dots,g_{-n})$ $=g_{1}c(\kern 0.8ptg_{1}^{-1}g_{2},\kern 0.8ptg_{2}^{-1}g_{3},\kern 0.8pt\dots,\kern 0.8ptg_{-n-1}^{-1}g_{-n}\kern 0.8pt),\quad\ n<0.$
We give formulas for cup product of cochains, homogeneous and inhomogeneous. The formulas for homogeneous cochains follow immediately from [3, Section IV.7, formulas for $\varphi_{p,q}$ on page 107]. We deduce formulas for inhomogeneous cochains from these formulas using formulas (4)–(7). Details of calculations are given in Appendix A.
Let AA and $B$ be two $G$-modules. For $m,n\in{\mathds{Z}}$, consider homogeneous cochains
$f^{\prime}_{m}\in F^{m}(G,A),\quad f^{\prime\prime}_{n}\in F^{n}(G,B),\quad f_{m+n}=f^{\prime}_{m}\cup f^{\prime\prime}_{n}\in F^{m+n}(G,A\otimes B).$
Consider the corresponding inhomogeneous cochains
$c^{\prime}_{m}\in C^{m}(G,A),\quad c^{\prime\prime}_{n}\in C^{n}(G,B).$
We find formulas for the inhomogeneous cochain $c_{m+n}\in C^{m+n}(G,A\otimes B)$ corresponding to $f_{m+n}$. We write then
$c_{m+n}=c^{\prime}_{m}\cup c^{\prime\prime}_{n}\kern 0.8pt.$
Let $p,q\geq 0$. We take $m=p$, $n=q$. Then we have
(8) $f_{p+q}$ $(g_{0},\dots,g_{p+q})=f^{\prime}_{p}(g_{0},\dots,g_{p})\,\otimes\,f^{\prime\prime}_{q}(g_{p},\dots,g_{p+q})$
(9) $c_{p+q}$ $(h_{1},\dots,h_{p+q})=$
$=c^{\prime}_{p}(h_{1},\dots,h_{p})\,\otimes\,(h_{1}h_{2}\cdots h_{p})\cdot c^{\prime\prime}_{q}(h_{p+1},\kern 0.8pt\dots,\kern 0.8pth_{p+q}).$
Let $p,q\geq 1$. We take m=pm=-p, n=qn=-q. Then we have
(10) $f_{-p-q}$ $(g_{0},\dots,g_{p+q-1})=f^{\prime}_{-p}(g_{0},\dots,g_{p-1})\,\otimes\,f^{\prime\prime}_{-q}(g_{p},\dots,g_{p+q-1}),$
(11) $c_{-p-q}$ $(h_{1},\dots,h_{p+q-1})=$
$=c^{\prime}_{-p}(h_{1},\dots,h_{p-1})\kern 0.8pt\otimes\kern 0.8pt(h_{1}h_{2}\cdots h_{p})\cdot c^{\prime\prime}_{-q}(h_{p+1},\kern 0.8pt\dots,\kern 0.8pth_{p+q-1}).$
Let $p\geq 0$, $q\geq 1$. We take $m=p$, n=pqn=-p-q. Then we have
(12) $f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\sum_{s}f^{\prime}_{p}(g_{1},\,s_{1},\ \dots,\ s_{p}\,)\,\otimes\,f^{\prime\prime}_{-p-q}(s_{p},\ \dots,\ s_{1},g_{1},\ \dots,\ g_{q}\,)$
(13) $c_{-q}(h_{1},\ \dots,\ h_{q-1})=$
$\quad=\sum_{t}c^{\prime}_{p}(\,t_{1},\ \dots,\ t_{p}\,)\,\otimes\,(t_{1}\cdots t_{p})\cdot c^{\prime\prime}_{-p-q}(\,t_{p}^{-1},\ \dots,\ t_{1}^{-1},h_{1},\ \dots,\ h_{q-1}\,).$
Let $p\geq 0$, $q\geq 1$. We take m=pqm=-p-q, $n=p$. Then we have
(14) $f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\sum_{s}f^{\prime}_{-p-q}(g_{1},\ \dots,\ g_{q},s_{1},\ \dots,\ s_{p})\,\otimes\,f^{\prime\prime}_{p}(s_{p},\ \dots,\ s_{1},g_{q}).$
(15) $c_{-q}(h_{1},\ \dots,\ h_{q-1})=$
$\ \ =\sum_{t}c^{\prime}_{-p-q}(h_{1},\ \dots,\ h_{q-1},t_{1},\ \dots,\ t_{p})\,\otimes\,(h_{1}h_{2}\cdots h_{q-1}t_{1}\cdots t_{p})\cdot c^{\prime\prime}_{p}(t_{p}^{-1},\ \dots,\ t_{1}^{-1}).$
Let $p\geq 0$, $q\geq 1$. We take $m=p+q$, n=qn=-q. Then we have
(16) $f_{p}(g_{0},\dots,g_{p})=\sum_{s}f^{\prime}_{p+q}(g_{0},\dots,g_{p},s_{1},\dots,s_{q})\,\otimes\,f^{\prime\prime}_{-q}(s_{q},\dots,s_{1})$
(17) $c_{p}(h_{1},\dots,h_{p})=$
$\quad=\sum_{t}c^{\prime}_{p+q}\big(\kern 0.8pth_{1},\dots,h_{p},t_{1},\dots,t_{q})\,\otimes\,(h_{1}\cdots h_{p}t_{1}\cdots t_{q})\cdot c^{\prime\prime}_{-q}(t_{q}^{-1},\dots,t_{2}^{-1})$
Let $p\geq 0$, $q\geq 1$. We take m=qm=-q, $n=p+q$. Then we have
(18) $f_{p}(g_{0},\ \dots,\ g_{p})=\sum_{s}f^{\prime}_{-q}(s_{1},\ \dots,\ s_{q})\,\otimes\,f^{\prime\prime}_{p+q}(s_{q},\ \dots,\ s_{1},g_{0},\ \dots,\ g_{p})$
(19) $c_{p}(h_{1},\ \dots,\ h_{p})=$
$\quad=\sum_{t}t_{1}\cdot c^{\prime}_{-q}(t_{2},\ \dots,\ t_{q})\,\otimes\,(t_{1}\cdots t_{q})\cdot c^{\prime\prime}_{p+q}(t_{q}^{-1},\ \dots,\ t_{1}^{-1},\ h_{1},\ \dots,\ h_{p}).$
In these formulas, the sums are taken over $s_{i}$ (resp., $t_{i}$) running independently through $G$.
3. The fundamental class of a cyclic extension
In this section we prove Theorem $1.$ Let $L/K$ be a finite Galois extension of non-archimedean local fields with cyclic Galois group $G={\rm Gal}(L/K)$ of order $n$. Choose a generator $\sigma\in G$. Let $e_{\sigma}\in K^{\times}$ be as in Subsection 1.2. We write $e_{\sigma}]$ for the image of $e_{\sigma}$ in $H^{-1}(G,L^{\times})\coloneqq K^{\times}/{\rm Nm}(L^{\times})$.
Consider the homomorphism
$\chi\colon G\to\frac{1}{n}{\mathds{Z}}/{\mathds{Z}}\quad\ \text{such that}\ \,\chi(\sigma)=\frac{1}{n}.$
Then $\chi\in H^{1}(G,{\mathds{Q}}/{\mathds{Z}})$. Consider
$\delta\chi\in H^{2}(G,{\mathds{Z}}),$
the image of $\chi$ under the connecting map
$\delta\colon H^{1}(G,\frac{1}{n}{\mathds{Z}}/{\mathds{Z}})\to H^{2}(G,{\mathds{Z}}).$
Then $\delta\chi$ is represented by the 2-cocycle
$b\in Z^{2}(G,{\mathds{Z}}),\quad\ b\colon G\times G\to{\mathds{Z}}$
defined as follows (compare Milne [5, page 102, above Proposition III.1.9]). We represent an element of $G\times G$ as $(\sigma^{i},\sigma^{j})$ with $i,j\in{\mathds{Z}}$, $0\leq i,j<n$. Then
$b(\sigma^{i},\sigma^{j})=\begin{cases}0&\text{if}\quad i+j<n\\ 1&\text{if}\quad i+j\geq n.\end{cases}$
For $a$ as in Theorem 1.1, we have
$a=b\cup e_{\sigma}$
where   $e_{\sigma}\in Z^{0}(G,L^{\times})=K^{\times}$   and $\cup$ denotes cup product of cocycles. It follows that $a$ is a 2-cocycle, $a\in Z^{2}(G,L^{\times})$. Write $a]\in H^{2}(G,L^{\times})$ for the class of $a$. Write
$e_{\sigma}]=e_{\sigma}\cdot{\rm Nm}(L^{\times})\in K^{\times}/{\rm Nm}(L^{\times})=\hat{H}^{0}(G,L^{\times}).$
Then
$a]=[b]\cup[e_{\sigma}]=\delta\chi\cup[e_{\sigma}]$
where $\cup$ denotes cup product of Tate cohomology classes. Consider the canonical isomorphism
${\rm inv}_{L/K}\colon H^{2}(G,L^{\times}){\ \xrightarrow{\makebox[9.56667pt]{\scriptsize${\raisebox{-3.01389pt}{$\sim$}}$}}\ }\frac{1}{n}{\mathds{Z}}/{\mathds{Z}}.$
We have
${\rm inv}_{L/K}(\delta\chi\cup[e_{\sigma}])=\chi(\sigma);$
see Harari [4, Proposition 9.3]. By construction, we have $\chi(\sigma)=\frac{1}{n}$. Thus
${\rm inv}_{L/K}[a]={\rm inv}_{L/K}(\delta\chi\cup[e_{\sigma}])=\chi(\sigma)=\frac{1}{n}.$
By the definition of the fundamental class $u_{L/K}$, it follows that $a]=u_{L/K}$, which completes the proof of Theorem $1.$
4. Explicit cocycles for a torus
In this section we prove Theorem 1.2. We sometimes write $z$ for $z_{x}$. We compute
$z=x\cup a\in Z^{-1}(G,X)\otimes Z^{2}(G,L^{\times})$
where
$z\in Z^{1}(G,L^{\times})\subseteq{\rm Maps}(G,L^{\times}),$
$a\in Z^{2}(G,L^{\times})\subseteq{\rm Maps}(G\times G,L^{\times}),$
$x\in Z^{-1}(G,X)\subseteq X.$
Here $a]=u_{L/K}$, the fundamental class, and the 2-cocycle $a$ is as in Theorem 1.1 and its proof, namely, $a=b\cup e_{\sigma}$.
Formula (19) for $p=1$, $q=1$ reads:
$z(g)=\sum_{t\in G}t\cdot x\,\otimes\,t\cdot a(t^{-1},g)$
for $g\in G$. Since $a$ takes the values $e_{\sigma}\in K^{\times}$ and 11 only, we have $t\cdot a(t^{-1},g)=a(t^{-1},g)$. It remains to compute the sum
$\sum_{t\in G}t\cdot x\,\otimes\,a(t^{-1},g).$
Recall that $\sigma$ is a generator of the cyclic group $G$. By abuse of notation, we write $\sigma^{g}$ for $g$ and $\sigma^{t}$ for $t$. Now $g,t\in{\mathds{Z}}/n{\mathds{Z}}$.
We wish to compute
$z(\sigma^{g})=\sum_{t\in{\mathds{Z}}/n{\mathds{Z}}}\sigma^{t}\cdot x\,\otimes\,a(\sigma^{-t},\sigma^{g}).$
For this end, we compute
$a(\sigma^{-t},\sigma^{g})=a(\sigma^{n-t},\sigma^{g})\quad\ \text{where}\ \,0\leq g<n,\quad 0<t\leq n.$
If $0<t\leq g<n$, then
$0<n-t<n,\quad 0\leq g<n,\quad(n-t)+g=n+(g-t)\geq n\ \ \text{(because $t\leq g$)},$
whence $a(\sigma^{n-t},\sigma^{g})=e_{\sigma}$.
If $t>g\geq 0,\ \,t\leq n$, then
$0\leq n-t<n,\quad 0\leq g<n,\quad(n-t)+g=n+(g-t)<n\ \ \text{(because $t>g$)},$
whence $a(\sigma^{n-t},\sigma^{g})=1$.
We conclude that
(20) $z_{x}(\sigma^{g})=\sum_{t=1}^{g}\sigma^{t}\cdot x\,\otimes\,e_{\sigma}$
for $0\leq g<n$. This completes the proof of Theorem 1.2.
Appendix A Cup product: calculations
A.1.
For $p,q\geq 0$, set $m=p$, $n=q$. Then by [3, Section IV.7, page 107] we have
(8) $f_{p+q}$ $(g_{0},\dots,g_{p+q})=f^{\prime}_{p}(g_{0},\dots,g_{p})\,\otimes\,f^{\prime\prime}_{q}(g_{p},\dots,g_{p+q}).$
Substituting (5) in the right-hand side, we obtain
$f_{p+q}(g_{0},\dots,g_{p+q})=g_{0}\cdot c^{\prime}_{p}(\kern 0.8ptg_{0}^{-1}g_{1},\kern 0.8pt\dots,\kern 0.8ptg_{p-1}^{-1}\kern 0.8ptg_{p}\kern 0.8pt)\,\otimes\,g_{p}\cdot c^{\prime\prime}_{q}(\kern 0.8ptg_{p}^{-1}g_{p+1},\kern 0.8pt\dots,\kern 0.8ptg_{p+q-1}^{-1}\kern 0.8ptg_{p+q}\kern 0.8pt).$
Substituting
$g_{0}=1,\ \,g_{1}=h_{1},\ \,g_{2}=h_{1}h_{2},\ \dots,\ g_{p}=h_{1}h_{2}\cdots h_{p},\ \dots,\ g_{p+q}=h_{1}h_{2}\cdots h_{p+q}\kern 0.8pt,$
we obtain
$g_{0}^{-1}g_{1}=h_{1},\ \,g_{1}^{-1}g_{2}=h_{2},\ \dots,\ g_{p-1}^{-1}\kern 0.8ptg_{p}=h_{p},$
$g_{p}^{-1}g_{p+1}=h_{p+1},\ \dots,\ g_{p+q-1}^{-1}\kern 0.8ptg_{p+q}=h_{p+q},$
whence
$f_{p+q}(1,h_{1},h_{1}h_{2},\ \dots,\ h_{1}h_{2}\cdots h_{p+q})=c^{\prime}_{p}(h_{1},\dots,h_{p})\kern 0.8pt\otimes\kern 0.8pt(h_{1}h_{2}\cdots h_{p})\cdot c^{\prime\prime}_{q}(h_{p+1},\kern 0.8pt\dots,\kern 0.8pth_{p+q}),$
and applying (4) to the left-hand side, we obtain
(9) $c_{p+q}(h_{1},\dots,h_{p+q})=c^{\prime}_{p}(h_{1},\dots,h_{p})\kern 0.8pt\otimes\kern 0.8pt(h_{1}h_{2}\cdots h_{p})\cdot c^{\prime\prime}_{q}(h_{p+1},\kern 0.8pt\dots,\kern 0.8pth_{p+q}).$
A.2.
For $p,q\geq 1$, set m=pm=-p, n=qn=-q. Then by [3, Section IV.7, page 107] we have
(10) $f_{-p-q}$ $(g_{0},\dots,g_{p+q-1})=f^{\prime}_{-p}(g_{0},\dots,g_{p-1})\,\otimes\,f^{\prime\prime}_{-q}(g_{p},\dots,g_{p+q-1}).$
Substituting (7) in the right-hand side, we obtain
$f_{-p-q}(g_{0},\dots,g_{p+q-1})=\\ =g_{0}\cdot c^{\prime}_{-p}(g_{0}^{-1}g_{1},g_{1}^{-1}g_{2},\dots,g_{p-2}^{-1}g_{p-1})\,\otimes\,g_{p}\cdot c^{\prime\prime}_{-q}(g_{p}^{-1}g_{p+1},\dots,g_{p+q-2}^{-1}g_{p+q-1}).$
Substituting
$g_{0}=1,\ \,g_{1}=h_{1},\ \,g_{2}=h_{1}h_{2},\ \dots,\ g_{p}=h_{1}h_{2}\cdots h_{p},\ \dots,\ g_{p+q-1}=h_{1}h_{2}\cdots h_{p+q-1}\kern 0.8pt,$
we obtain
$g_{0}^{-1}g_{1}=h_{1},\ \,g_{1}^{-1}g_{2}=h_{2},\ \dots,\ g_{p-2}^{-1}\kern 0.8ptg_{p-1}=h_{p-1},$
$g_{p}^{-1}g_{p+1}=h_{p+1},\ \dots,\ g_{p+q-2}^{-1}\kern 0.8ptg_{p+q-1}=h_{p+q-1},$
whence
$f_{-p-q}(1,h_{1},h_{1}h_{2},\dots,h_{1}h_{2}\cdots h_{p+q-1})=\\ =c^{\prime}_{-p}(h_{1},\dots,h_{p-1})\kern 0.8pt\otimes\kern 0.8pt(h_{1}h_{2}\cdots h_{p})\cdot c^{\prime\prime}_{-q}(h_{p+1},\kern 0.8pt\dots,\kern 0.8pth_{p+q-1}),$
and applying (6) to the left-hand side, we obtain
(11) $c_{-p-q}(h_{1},\dots,h_{p+q-1})=c^{\prime}_{-p}(h_{1},\dots,h_{p-1})\kern 0.8pt\otimes\kern 0.8pt(h_{1}h_{2}\cdots h_{p})\cdot c^{\prime\prime}_{-q}(h_{p+1},\kern 0.8pt\dots,\kern 0.8pth_{p+q-1}).$
A.3.
From now on we assume that $p\geq 0$, $q\geq 1$.
We take $m=p$, n=pqn=-p-q. Then by [3, Section IV.7, page 107] we have
(12) $f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\sum_{s}f^{\prime}_{p}(g_{1},\,s_{1},\ \dots,\ s_{p}\,)\,\otimes\,f^{\prime\prime}_{-p-q}(s_{p},\ \dots,\ s_{1},g_{1},\ \dots,\ g_{q}\,)$
where $s_{i}$ run independently through $G$. Substituting (5) and (7) in the right-hand side, we obtain
$f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\sum_{s}\Big(g_{1}\cdot c^{\prime}_{p}(g_{1}^{-1}s_{1},s_{1}^{-1}s_{2},\ \dots,\ s_{p-1}^{-1}s_{p})\\ \,\otimes\,s_{p}\cdot c^{\prime\prime}_{-p-q}(s_{p}^{-1}s_{p-1},\,\dots,\,s_{2}^{-1}s_{1},s_{1}^{-1}g_{1},g_{1}^{-1}g_{2},\,\dots,\,g_{q-1}^{-1}g_{q}\Big).$
Substituting
$s_{1}=t_{1},\,s_{2}=t_{1}t_{2},\ \dots,\ s_{p}=t_{1}\cdots t_{p},$
$g_{1}=1,\ g_{2}=h_{1},\,g_{3}=h_{1}h_{2},\ \dots,\ g_{q}=h_{1}\cdots h_{q-1},$
we obtain
$g_{1}^{-1}s_{1}=t_{1},\ s_{1}^{-1}s_{2}=t_{2},\ \dots,\ s_{p-1}^{-1}s_{p}=t_{p},$
$s_{p}^{-1}s_{p-1}=t_{p}^{-1},\ \dots,\ s_{2}^{-1}s_{1}=t_{2}^{-1},\ s_{1}^{-1}g_{1}=t_{1}^{-1},$
$g_{1}^{-1}g_{2}=h_{1},\ g_{2}^{-1}g_{3}=h_{2},\ \dots,\ g_{q-1}^{-1}g_{q}=h_{q-1},$
whence
$f_{-q}(\,1,h_{1},h_{1}h_{2},\ \dots,\ h_{1}h_{2}\cdots h_{q-1}\,)=\\ =\sum_{t}c^{\prime}_{p}(\,t_{1},\ \dots,\ t_{p}\,)\,\otimes\,(t_{1}\cdots t_{p})\cdot c^{\prime\prime}_{-p-q}(\,t_{p}^{-1},\ \dots,\ t_{1}^{-1},h_{1},\ \dots,\ h_{q-1}\,),$
and applying (6) to the left-hand side, we obtain
(13) $c_{-q}(h_{1},\ \dots,\ h_{q-1})=\sum_{t}c^{\prime}_{p}(\,t_{1},\ \dots,\ t_{p}\,)\,\otimes\,(t_{1}\cdots t_{p})\cdot c^{\prime\prime}_{-p-q}(\,t_{p}^{-1},\ \dots,\ t_{1}^{-1},h_{1},\ \dots,\ h_{q-1}\,)$
where $t_{i}$ run independently through $G$.
A.4.
We take m=pqm=-p-q, $n=p$. Then by [3, Section IV.7, page 107] we have
(14) $f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\sum_{s}f^{\prime}_{-p-q}(g_{1},\ \dots,\ g_{q},s_{1},\ \dots,\ s_{p})\,\otimes\,f^{\prime\prime}_{p}(s_{p},\ \dots,\ s_{1},g_{q})$
where $s_{i}$ run independently through $G$. Substituting (7) and (5) in the right-hand side, we obtain
$f_{-q}(\,g_{1},\ \dots,\ g_{q}\,)=\\ =\sum_{s}g_{1}\cdot c^{\prime}_{-p-q}(g_{1}^{-1}g_{2},\ \dots,\ g_{q-2}^{-1}g_{q-1},\ g_{q}^{-1}s_{1},\ s_{1}^{-1}s_{2},\ \dots,s_{p-1}^{-1}s_{p})\\ \otimes\,s_{p}\cdot c^{\prime\prime}_{p}(s_{p}^{-1}s_{p-1},\ \dots,\ s_{2}^{-1}s_{1},s_{1}^{-1}g_{q}).$
Substituting
$g_{1}=1,\ g_{2}=h_{1},\ g_{3}=h_{1}h_{2},\ \dots,\ g_{q}=h_{1}h_{2}\cdots h_{q-1},$
$s_{1}=h_{1}h_{2}\cdots h_{q-1}t_{1},\ \dots,\ s_{p}=h_{1}h_{2}\cdots h_{q-1}t_{1}\cdots t_{p}\kern 0.8pt,$
we obtain
$g_{1}^{-1}g_{2}=h_{1},\ \dots,\ g_{q-1}^{-1}g_{q}=h_{q-1},$
$g_{q}^{-1}s_{1}=t_{1},\ s_{1}^{-1}s_{2}=t_{2},\ \dots,\ s_{p-1}^{-1}s_{p}=t_{p},$
$s_{p}^{-1}s_{p-1}=t_{p}^{-1},\ \dots,\ s_{2}^{-1}s_{1}=t_{2}^{-1},\ s_{1}^{-1}g_{q}=t_{1}^{-1},$
whence
$f_{-q}(\,1,\ h_{1},\ h_{1}h_{2},\ \dots,\ h_{1}h_{2}\cdots h_{q-1}\,)=\\ =\sum_{t}c^{\prime}_{-p-q}(h_{1},\ \dots,\ h_{q-1},t_{1},\ \dots,\ t_{p})\,\otimes\,(h_{1}h_{2}\cdots h_{q-1}t_{1}\cdots t_{p})\cdot c^{\prime\prime}_{p}(t_{p}^{-1},\ \dots,\ t_{1}^{-1}),$
and applying (6) to the left-hand side, we obtain
$c_{-q}(\,h_{1},\ \dots,\ h_{q-1}\,)=\\ \sum_{t}c^{\prime}_{-p-q}(h_{1},\ \dots,\ h_{q-1},t_{1},\ \dots,\ t_{p})\,\otimes\,(h_{1}h_{2}\cdots h_{q-1}t_{1}\cdots t_{p})\cdot c^{\prime\prime}_{p}(t_{p}^{-1},\ \dots,\ t_{1}^{-1})$
where $t_{i}$ run independently through $G$.
A.5.
We take $m=p+q$, n=qn=-q. Then by [3, Section IV.7, page 107] we have
(16) $f_{p}(g_{0},\dots,g_{p})=\sum_{s}f^{\prime}_{p+q}(g_{0},\dots,g_{p},s_{1},\dots,s_{q})\,\otimes\,f^{\prime\prime}_{-q}(s_{q},\dots,s_{1})$
where $s_{i}$ run independently through $G$. Substituting (5) and (7) in the right-hand side, we obtain
$f_{p}(\,g_{0},\ \dots,\ g_{p}\,)=\\ =\sum_{s}\Big(g_{0}\cdot c^{\prime}_{p+q}\big(\kern 0.8ptg_{0}^{-1}g_{1},\ \dots,\ g_{p-1}^{-1}g_{p},\ g_{p}^{-1}s_{1},\ s_{1}^{-1}s_{2},\dots,\ s_{q-1}^{-1}s_{q}\kern 0.8pt\big)\\ \,\otimes\,s_{q}\cdot c^{\prime\prime}_{-q}(s_{q}^{-1}s_{q-1},\ \dots,\ s_{2}^{-1}s_{1})\Big).$
Substituting
$g_{0}=1,\ \,g_{1}=h_{1},\ \,g_{2}=h_{1}h_{1},\ \dots,\ g_{p}=h_{1}h_{2}\cdots h_{p},$
$s_{1}=h_{1}h_{2}\cdots h_{p}t_{1},\ s_{2}=h_{1}h_{2}\cdots h_{p}t_{1}t_{2},\ \dots,\ s_{q}=h_{1}h_{2}\cdots h_{p}t_{1}\cdots t_{q}\kern 0.8pt,$
we obtain
$g_{0}^{-1}g_{1}=h_{1},\ \,g_{1}^{-1}g_{2}=h_{2},\ \dots,\ \,g_{p-1}^{-1}g_{p}=h_{p},$
$g_{p}^{-1}s_{1}=t_{1},\ \,s_{1}^{-1}s_{2}=t_{2},\ \dots,\ s_{q-1}^{-1}s_{q}=t_{q},$
$s_{q}^{-1}s_{q-1}=t_{q}^{-1},\ \dots,\ s_{2}^{-1}s_{1}=t_{2}^{-1},$
whence
$f_{p}(1,h_{1},h_{1}h_{2},\ \dots,\ h_{1}h_{2}\cdots h_{p})=\\ \sum_{t}c^{\prime}_{p+q}\big(\,\kern 0.8pth_{1},\ \dots,\ h_{p},t_{1},\ \dots,\ t_{q}\,\big)\,\otimes\,(h_{1}\cdots h_{p}t_{1}\cdots t_{q})\cdot c^{\prime\prime}_{-q}(t_{q}^{-1},\ \dots,\ t_{2}^{-1}),$
and applying (4) to the left-hand side, we obtain
$c_{p}(h_{1},\dots,h_{p})=\sum_{t}c^{\prime}_{p+q}\big(\,\kern 0.8pth_{1},\ \dots,\ h_{p},t_{1},\ \dots,\ t_{q}\,\big)\,\otimes\,(h_{1}\cdots h_{p}t_{1}\cdots t_{q})\cdot c^{\prime\prime}_{-q}(t_{q}^{-1},\ \dots,\ t_{2}^{-1})$
where $t_{i}$ run independently through $G$.
A.6.
We take m=qm=-q, $n=p+q$. Then by [3, Section IV.7, page 107] we have
(18) $f_{p}(g_{0},\ \dots,\ g_{p})=\sum_{s}f^{\prime}_{-q}(s_{1},\ \dots,\ s_{q})\,\otimes\,f^{\prime\prime}_{p+q}(s_{q},\ \dots,\ s_{1},g_{0},\ \dots,\ g_{p})$
where $s_{i}$ run independently through $G$. Substituting (7) and (5) in the right-hand side, we obtain
$f_{p}(g_{0},\ \dots,\ g_{p})=\\ =\sum_{s}s_{1}\cdot c^{\prime}_{-q}(s_{1}^{-1}s_{2},\ \dots,\ s_{q-1}^{-1}s_{q})\,\otimes\,s_{q}\cdot c^{\prime\prime}_{p+q}(s_{q}^{-1}s_{q-1},\ \dots,\ s_{1}^{-1}g_{0},\ g_{0}^{-1}g_{1},\ \dots,\ g_{p-1}^{-1}g_{p}).$
Substituting
$g_{0}=1,\ \,g_{1}=h_{1},\ \,g_{2}=h_{1}h_{1},\ \dots,\ g_{p}=h_{1}h_{2}\cdots h_{p},$
$s_{1}=t_{1},s_{2}=t_{1}t_{2},\ \dots,\ s_{q}=t_{1}\cdots t_{q}\kern 0.8pt,$
we obtain
$s_{1}^{-1}s_{2}=t_{2},\ \dots,\ s_{q-1}^{-1}s_{q}=t_{q},$
$s_{q}^{-1}s_{q-1}=t_{q}^{-1},\ \dots,\ s_{1}^{-1}g_{0}=t_{1}^{-1}\kern 0.8pt,$
$g_{0}^{-1}g_{1}=h_{1},\ \dots,\ g_{p-1}^{-1}g_{p}=h_{p},$
whence
$f_{p}(1,\ h_{1},\ \dots,\ h_{1}\cdots h_{p})=\\ =\sum_{t}t_{1}\cdot c^{\prime}_{-q}(t_{2},\ \dots,\ t_{q})\,\otimes\,(t_{1}\cdots t_{q})\cdot c^{\prime\prime}_{p+q}(t_{q}^{-1},\ \dots,\ t_{1}^{-1},\ h_{1},\ \dots,\ h_{p}),$
and applying (4) to the left-hand side, we obtain
(19) $c_{p}(h_{1},\ \dots,\ h_{p})=\sum_{t}t_{1}\cdot c^{\prime}_{-q}(t_{2},\ \dots,\ t_{q})\,\otimes\,(t_{1}\cdots t_{q})\cdot c^{\prime\prime}_{p+q}(t_{q}^{-1},\ \dots,\ t_{1}^{-1},\ h_{1},\ \dots,\ h_{p})$
where $t_{i}$ run independently through $G$.